OpenStudy (anonymous):

A student increases the temperature of a 556 cm3 balloon from 278 K to 308 K. Assuming constant pressure, what should the new volume of the view be? 320 cm3 501 cm3 1232 cm3 616 cm3

OpenStudy (stephstar101):

P1*V1=P2*V2 - equation (278 K)(556 cm3)=(308 K)*V - plug in values 154568 Kcm3=(308 K)*V - multiply the left side V= 501 cm3 - divide both sides by 308K and you get your answer